the longer leg of a right triangle is 1 inch longer than the shorter leg. the hypotenuse is 9 inches longer than the shorter leg. find the side lengths of the triangle.
Answer by stanbon(75887)

the longer leg of a right triangle is 1 inch longer than the shorter leg. the hypotenuse is 9 inches longer than the shorter leg. find the side lengths of the triangle.
------------------
shorter leg:: x
longer leg:: x+1
hypotenuse:: x+9
------------------------
x^2 + (x+1)^2 = (x+9)^2
-----
x^2 + x^2+2x+1 = x^2 + 18x + 81
--------
x^2-16x-80 = 0
(x-20)(x+4) = 0
Positive solution:
x = 20 in (shorter)
x+1 = 21 in (longer)
x+9 = 29 in (hypotenuse)
-------------
Cheers,
Stan H.
--------------